OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation x^4-41x^2=-400

OpenStudy (ddcamp):

Let u=x² Then solve: u²-41u+400 = 0 The x values will be ±sqrt(u)

OpenStudy (anonymous):

this is an equation that can be reduced to a quadratic equation \[x ^{4}-41x ^{2}+400=0\] so you have factors of 400 that add to -41 are -25 and - 16 so you need to solve \[(x ^{2}−16)(x ^{2}−25)=0\] each quadratic in the brackets above are the difference of 2 squares. Hope this helps

OpenStudy (anonymous):

what do you mean about the difference of 2squares

OpenStudy (ddcamp):

The difference of two squares rule: (a²-b²) = (a-b)(a+b) Since 16 = 4², we can say: (x² - 16) = (x+4)(x-4)